# Hierarchical Neural Simulation-Based Inference Over Event Ensembles

### v1

@@ 3.2 Mixture models in particle physics: frequentist treatment
We demonstrate the dataset level profiling by adapting the approach described in Heinrich [13] and compare the trained test statistic to the true profile likelihood ratio.
The forward model:
@@
with $c_{s}=c_{s}(\theta,\theta_{\nu})$ being the relative signal strength and $p_{s}$ and $p_{b}$ the per-event densities for signal and background, respectively. Fast inference for such models is especially important in settings that perform real-time monitoring of the quality of the data stream delivered by the experiment. The observed per-event densities are typically not tractable, but high-fidelity simulators exist. In this work we focus on Gaussian mixture components, since the empirical distributions typically considered in the domain use-cases feature multi-modal densities (e.g., in searches for “bump” features).
To demonstrate the simulation-based frequentist inference, we consider the special case where $p_{s}$ and $p_{b}$ are Normal distributions $\mathrm{No}(\mu,\sigma)$ with fixed hyperparameters $\mu_{s/b},\sigma_{s/b}$, but the component coefficients are a function of both the parameter of interest (the “signal strength”) $\theta$ as well as the nuisance parameter (i.e. “the background strength”): $c_{s}=\theta/(\theta+\theta_{\nu})$. The overall expected rate is then $\lambda(\theta,\theta_{\nu})=\theta+\theta_{\nu}$.
Inference:
The model is again based on the deep set-inspired architecture described in Sec. 2.2, with a small variation – after aggregating the per-event embeddings, these are concatenated with the parameters of interest $\theta$ in order to obtain a parameterized neural network classifier one-to-one with the desired test statistic [6], $s_{\phi}(\{x\},\theta)\overset{\sim}{\leftrightarrow}t_{\theta}(\{x\})$. Following the procedure in [13] we aim to train a test statistic with best average power across alternatives irrespective of the nuisance parameters and show that the result is a good approximation of the profile likelihood ratio, i.e., the dataset-wide test statistic typically used for interval estimation and hypothesis testing in frequentist inference:
@@
Results:
@@ 3.3 Mixture models in particle physics: Bayesian inference for a narrow resonance
The forward model:
We also demonstrate a Bayesian dataset-wide inference for the mixture model case. Here the parameters of the model affect the mean of the signal component and the mixture coefficient, whereas the signal width and all background parameters are fixed hyperparameters. Using the notation in Eq. (8), we have
@@
i.e. the parameter of interest $\theta$ is the relative signal strength and the location of the signal in the observed feature space is a nuisance parameter. In typical applications in particle physics, the signal is often very narrow, whereas the background distribution is more diffuse. All parameter choices and the priors $p(\theta)$ and $p(\theta_{\nu})$ are specified in the App. A.3. Instead of parameterizing the overall Poisson rate $\lambda$, here we used a fixed batch of events of size $N_{0}$, corresponding to the situation one might encounter in a real-time inference task. In summary, the dataset-wide joint probability is:
@@
Figure 4: Signal- and background densities $p_{s}$ and $p_{b}$ in a simplified particle physics analysis for two different true values of the nuisance parameter $\theta_{\nu}$ (left), additionally showing sample datasets $\{x\}$ of size $N_{0}=100$. The median mean $\mu_{\th\theta}$ (center) and median standard deviation $\sigma_{\theta}$ (right) of the posterior $p(\theta\mid\{x\})$, obtained with different inference methods for different true nuisance parameter values $\theta_{\nu,\mathrm{true}}$. The median is computed on ensembles of 400 datasets for each parameter point.
We use a deep set network to perform fast, amortized inference in this setting, trained as described in Sec. 2.2 directly on the set of observations $\{x\}$. Using the likelihood of Eq. (11), the posterior is also accessible through MCMC inference (as described in the App. A.3), which can serve as a point of comparison.
In many cases, the individual events $x$ are too high-dimensional for direct inference to be possible. A fundamental problem in the analysis of collider data thus consists in the construction of a low-dimensional summary statistic $x\rightarrow s(x)$, designed to retain the relevant information contained in the data, subsequently used in the inference step.
It is common practice in contemporary work to use the signal-to-background density ratio $s_{\mathrm{nom}}(x;\theta_{\nu,\mathrm{nom}})=p_{s}(x\mid\theta_{\nu,\mathrm{nom}})/p_{b}(x)$ evaluated at a particular “nominal” nuisance parameter value $\theta_{\nu,\mathrm{nom}}$. This observable is known to be a sufficient statistic for $\theta$ under the hypothesis $\theta_{\nu}=\theta_{\nu,\mathrm{nom}}$ [19].
Another popular choice is to neglect the hierarchical structure of the data-generating process; in the Bayesian paradigm this involves marginalizing over the nuisance parameters at a per-event level and using $s_{\mathrm{marg}}(x)=\int\mathrm{d}\theta_{\nu}\,p_{s}(x\mid\theta_{\nu})\,p(\theta_{\nu})/p_{b}(x)$ as summary statistic. Both quantities are known to not be information-preserving in general. Our example admits analytic likelihoods also for $s_{\mathrm{nom}}$ and $s_{\mathrm{marg}}$, allowing their performance to be evaluated through the MCMC procedure described above.
Results and discussion
@@
Using the fully parameterized density ratio $p_{s}(x\mid\theta_{\nu})/p_{b}(x)$ for optimal inference is, as mentioned in Sec. 2.1, complicated by the fact that often many hundred nuisance parameters are present in contemporary particle physics problems [26, 27]. However, nuisance parameters inform the the deep set through their presence in the training dataset and do not require any explicit parameterization of the network, suggesting a much better scaling behavior.
@@ 3.4 Astrophysics example: Strong gravitational lensing
We test our procedure for hierarchical simulation-based inference on a more complicated problem from the domain of astrophysics, where the likelihood associated with the forward model is intractable – the characterization of the distribution of dark matter clumps (subhalos) in galaxies using images of background galaxies gravitationally lensed by the clumps as well as a larger smooth mass distribution containing them. The model contains both per-event (i.e., local) as well as dataset-wide (i.e., global) parameters, showcasing the ability to do dataset-wide hierarchical simulation-based inference.
The forward model: The substructure mass function, which describes the mass distribution of clumps, is parameterized as a power law $\frac{\mathrm{d}n}{\mathrm{d}m_{\mathrm{sub}}}=\alpha\cdot m_{\mathrm{sub}}^{-\beta}$ where the normalization $\alpha$ parameterizes the overall abundance of subhalos, and $\beta$ is the spectral index. $\theta_{\mathrm{glob}}=\{\alpha,\beta\}$ are the global parameters in the hierarchical model. The number of subhalos follows a Poisson rate, $N_{\mathrm{sub}}\sim\mathrm{Pois}(\mu_{\mathrm{sub}})$, where we have the expected number of subhalos $\mu_{\mathrm{sub}}=\int\mathrm{d}m_{\mathrm{sub}}\,\frac{\mathrm{d}n}{\mathrm{d}m_{\mathrm{sub}}}$. Heavier clumps which are closer to the lensed ring (Einstein ring) generally have a larger effect on the lensing image.
In addition we have per-image (local) parameters $\theta_{\mathrm{loc}}=\{\theta_{E},\Delta x,\Delta y,q\}$, where $\theta_{E}$ is the Einstein radius characterizing the overall scale of the lensing ring, $\{\Delta x,\Delta y\}$ is the offset vector between the centers of the background source and the lens galaxy, and $q\in[0,1]$ is the axis ratio with 1 corresponding to a spherical lens. Subhalos are uniformly placed within radial distance $0.2^{\prime\prime}$ of the Einstein radius $\theta_{E}$.
Given a realization of the masses and positions of the clumps $\{x_{\mathrm{sub}},y_{\mathrm{sub}},m_{\mathrm{sub}}\}_{i=1}^{N_{\mathrm{sub}}}$ (which otherwise act as latent variables, collectively denoted $\{z_{\mathrm{sub}}\}$), the expected image $\mu_{\mathrm{image}}$ can be obtained using a gravitational lensing simulator [4]. The actual image is then a Poisson realization of this expectation, $x\sim\mathrm{Pois}(\mu_{\mathrm{image}})$. We choose our simulated lensed images to be $64\times 64$ pixels in size; see the App. A.4 for more details on the lensing forward model and details on experiments.
The per-image likelihood is intractable because of the high dimensionality of the latent space, described by the number, masses, and positions of individual subhalo clumps.
@@
Inference: We would like to infer the posterior over the local as well as global parameters $p(\theta_{\mathrm{glob}},\theta_{\mathrm{loc}}\mid\{x\})$ given an ensemble of images. As discussed in Sec. 2, building per-image posterior estimators ignores the joint influence of the global parameters over the entire dataset, potentially reducing the identifiability of the local parameters. If a combined posterior for the global parameters is desired, combining per-image estimators into a global posterior can also be error-prone as per-observation errors accumulate, especially if flexible density estimators (e.g., normalizing flows) are used, as we do here.
@@
We train a hierarchical neural posterior estimator with datasets containing up to 25 lensed images, using a ResNet-18 [12] CNN as the per-event feature extractor. Figure 6 shows exemplary posterior inference results on one of the local parameters, the Einstein radius $\theta_{E}$ (left) and one of the global parameters, the amplitude of the clump mass function $\alpha$ (right), on test datasets. The red lines show posteriors using the hierarchical estimator, while the blue dotted lines show results of using a per-image estimator. We can see that the posterior mass concentrates around the true value (vertical dashed) as more images are analyzed. On the other hand, the local parameter values are very similar to those obtained without simultaneously constraining dataset-wide global parameters. This is because, in this case, sub-clumps have a specific effect on the lensing image, and the degeneracy between the local and global parameters is weak.
@@
Figure 6: Example results of dataset-wide inference on local (left) as well as global (right) gravitational lensing parameters using the hierarchical simulation-based inference model. Concentration of the posterior mass around the true global parameter value is seen as more lensed images are analyzed, and the local parameters are correctly recovered.
@@ 4 Conclusions
We studied the problem of simulation-based inference over event ensembles in the presence of hierarchical structures in the underlying forward model. In this setting, common in the physical sciences, optimal parameter inference requires a dataset-wide approach. For this purpose, we introduced neural estimators for likelihood-ratios and posteriors, illustrated their use for Bayesian as well as frequentist analysis in situations typical of astro- and particle physics, and substantiated optimal inference performance in these scenarios. Our methods do not require explicit parameterization in terms of marginalized or profiled nuisance parameters, making them complementary to existing techniques for dataset-wide inference. Once trained, they provide fully amortized inference over datasets of varying cardinality and offer significant speedups over traditional frequentist and Bayesian methods, making them suitable also for high-throughput applications.
Limitations: Our methods apply to situations where the cardinality $N$ of the observed dataset follows an arbitrary, but specified, distribution $p(N\mid\theta)$ (cf. Eq. (2)). It remains to be systematically explored in the hierarchical setting to what extent estimators trained on $p_{1}(N\mid\theta)$ can achieve statistically sound performance when applied to datasets corresponding to a different distribution $p_{2}(N\mid\theta)$. We leave this extension for future investigation, but note that insights from recent work describing Bayesian updating as a dynamical system [2, 5] or using score-based models [10] may offer promising ways to address this problem.
Acknowledgments and Disclosure of Funding
We made extensive use of the Einops [24], Jax [3], Jupyter [16], Matplotlib [14], nflows [8], Numpy [11], PyMC5 [25], PyTorch [21], PyTorch Lightning [9], and Scipy [29] packages.

### v2

@@ 5.2 Mixture models in particle physics: frequentist treatment
We demonstrate the dataset level profiling by adapting the approach described in Heinrich (2022) and compare the trained test statistic to the true profile likelihood ratio (not to be confused by the likelihood-to-evidence ratio (Miller et al., 2022)).
The forward model
@@
with $c_{s}=c_{s}(\theta,\theta_{\nu})$ being the relative signal strength and $p_{s}$ and $p_{b}$ the per-event densities for signal and background, respectively. Fast inference for such models is especially important in settings that perform real-time monitoring of the quality of the data stream delivered by the experiment. The observed per-event densities are typically not tractable, but high-fidelity simulators exist. In this work we focus on Gaussian mixture components, since the empirical distributions typically considered in the domain use-cases feature multi-modal densities (e.g., in searches for “bump” features). To demonstrate the simulation-based frequentist inference, we consider the special case where $p_{s}$ and $p_{b}$ are Normal distributions $\mathrm{No}(\mu,\sigma)$ with fixed hyperparameters $\mu_{s/b},\sigma_{s/b}$, but the component coefficients are a function of both the parameter of interest (the “signal strength”) $\theta$ as well as the nuisance parameter (i.e. “the background strength”): $c_{s}=\theta/(\theta+\theta_{\nu})$. The overall expected rate is then $\lambda(\theta,\theta_{\nu})=\theta+\theta_{\nu}$.
Inference
The model is again based on the deep set-inspired architecture described in Sec. 4, with a small variation – after aggregating the per-event embeddings, these are concatenated with the parameters of interest $\theta$ in order to obtain a parameterized neural network classifier one-to-one with the desired test statistic (Cranmer et al., 2015), $s_{\phi}(\{x\},\theta)\overset{\sim}{\leftrightarrow}t_{\theta}(\{x\})$. Following the procedure in Heinrich (2022) we aim to train a test statistic with best average power across alternatives irrespective of the nuisance parameters and show that the result is a good approximation of the profile likelihood ratio, i.e., the dataset-wide test statistic typically used for interval estimation and hypothesis testing in frequentist inference:
@@
Details on the training procedure as well as hyperparameter choices are provided in App. A.
Results
@@ 5.3 Mixture models in particle physics: Bayesian inference for a narrow resonance
The forward model
We also demonstrate a Bayesian dataset-wide inference for the mixture model case. Here the parameters of the model affect the mean of the signal component and the mixture coefficient, whereas the signal width and all background parameters are fixed hyperparameters. Using the notation in equation 8, we have
@@
i.e. the parameter of interest $\theta$ is the relative signal strength and the location of the signal in the observed feature space is a nuisance parameter. In typical applications in particle physics, the signal is often very narrow, whereas the background distribution is more diffuse. All parameter choices and the priors $p(\theta)$ and $p(\theta_{\nu})$ are specified in App. A. Instead of parameterizing the overall Poisson rate $\lambda$, here we used a fixed batch of events of size $N_{0}$, corresponding to the situation one might encounter in a real-time inference task. In summary, the dataset-wide joint probability is:
@@
Figure 4: Signal- and background densities $p_{s}$ and $p_{b}$ in a simplified particle physics analysis for two different true values of the nuisance parameter $\theta_{\nu}$ (left), additionally showing sample datasets $\{x\}$ of size $N_{0}=100$. The median mean $\mu_{\theta}$ (center) and median standard deviation $\sigma_{\theta}$ (right) of the posterior $p(\theta\mid\{x\})$, obtained with different inference methods for different true nuisance parameter values $\theta_{\nu,\mathrm{true}}$. The median is computed on ensembles of 400 datasets for each parameter point.
We use a deep set network to perform fast, amortized inference in this setting, trained as described in Sec. 4 directly on the set of observations $\{x\}$. Using the likelihood of equation 11, the posterior is also accessible through MCMC inference (as described in App. A), which can serve as a point of comparison. In many cases, the individual events $x$ are too high-dimensional for direct inference to be possible. A fundamental problem in the analysis of collider data thus consists in the construction of a low-dimensional summary statistic $x\rightarrow s(x)$, designed to retain the relevant information contained in the data, subsequently used in the inference step. It is common practice in contemporary work to use the signal-to-background density ratio $s_{\mathrm{nom}}(x;\theta_{\nu,\mathrm{nom}})=p_{s}(x\mid\theta_{\nu,\mathrm{nom}})/p_{b}(x)$ evaluated at a particular “nominal” nuisance parameter value $\theta_{\nu,\mathrm{nom}}$. This observable is known to be a sufficient statistic for $\theta$ under the hypothesis $\theta_{\nu}=\theta_{\nu,\mathrm{nom}}$ (Neyman & ES, 1933). Another popular choice is to neglect the hierarchical structure of the data-generating process; in the Bayesian paradigm this involves marginalizing over the nuisance parameters at a per-event level and using $s_{\mathrm{marg}}(x)=\int\mathrm{d}\theta_{\nu}\,p_{s}(x\mid\theta_{\nu})\,p(\theta_{\nu})/p_{b}(x)$ as summary statistic. Both quantities are known to not be information-preserving in general. Our example admits analytic likelihoods also for $s_{\mathrm{nom}}$ and $s_{\mathrm{marg}}$, allowing their performance to be evaluated through the MCMC procedure described above.
Results
@@
Using the fully parameterized density ratio $p_{s}(x\mid\theta_{\nu})/p_{b}(x)$ for hierarchy-aware inference is, as mentioned in Sec. 2.3, complicated by the fact that often many hundred nuisance parameters are present in contemporary particle physics problems (The ATLAS Collaboration, 2022; The CMS Collaboration, 2022). However, nuisance parameters inform the the deep set through their presence in the training dataset and do not require any explicit parameterization of the network, suggesting a much better scaling behavior.
@@ 5.4 Astrophysics example: Strong gravitational lensing
We test our procedure for hierarchical simulation-based inference on a more complicated problem from the domain of astrophysics, where the likelihood associated with the forward model is intractable – the characterization of the distribution of dark matter clumps (subhalos) in galaxies using images of galaxies gravitationally lensed by the clumps as well as a larger smooth mass distribution. The model contains both per-event (i.e., local) as well as dataset-wide (i.e., global) parameters. Estimating global lensing parameters with more sophisticated forward models was studied in, e.g., Anau Montel et al. (2023); Montel & Weniger (2022); Coogan et al. (2022); Wagner-Carena et al. (2022); Brehmer et al. (2019); we note that our aim here is instead to showcase the ability to do hierarchical simulation-based inference over local and global parameters with an amortized set-wide estimator.
The forward model
The substructure mass function, which describes the mass distribution of clumps, is parameterized as a power law $\frac{\mathrm{d}n}{\mathrm{d}m_{\mathrm{sub}}}=\alpha\cdot m_{\mathrm{sub}}^{-\beta}$ where the normalization $\alpha$ parameterizes the overall abundance of subhalos, and $\beta$ is the spectral index. $\theta_{\mathrm{glob}}=\{\alpha,\beta\}$ are the global parameters in the hierarchical model. The number of subhalos follows a Poisson rate, $N_{\mathrm{sub}}\sim\mathrm{Pois}(\mu_{\mathrm{sub}})$, where we have the expected number of subhalos $\mu_{\mathrm{sub}}=\int\mathrm{d}m_{\mathrm{sub}}\,\frac{\mathrm{d}n}{\mathrm{d}m_{\mathrm{sub}}}$. In addition we have per-image (local) parameters $\theta_{\mathrm{loc}}=\{\theta_{E},\Delta x,\Delta y,q\}$, where $\theta_{E}$ is the Einstein radius characterizing the overall scale of the lensing ring, $\{\Delta x,\Delta y\}$ is the offset vector between the centers of the background source and the lens galaxy, and $q\in[0,1]$ is the axis ratio with 1 corresponding to a spherical lens. The subhalo radial positions are drawn from a uniform distribution close to the Einstein radius, $\{r_{\mathrm{sub}}\}_{i=1}^{N_{\mathrm{sub}}}\sim\mathrm{Unif}(\theta_{E}-0.2^{\prime\prime},\theta_{E}+0.2^{\prime\prime})$ and the azimuthal angle is drawn uniformly from the interval $[0,2\pi]$. Given a realization of the masses and positions of the clumps $\{x_{\mathrm{sub}},y_{\mathrm{sub}},m_{\mathrm{sub}}\}_{i=1}^{N_{\mathrm{sub}}}$ (which otherwise act as latent variables, collectively denoted $\{z_{\mathrm{sub}}\}$), the expected image $\mu_{\mathrm{image}}$ can be obtained using a gravitational lensing simulator (Brehmer et al., 2019). The actual image is then a Poisson realization of this expectation, $x\sim\mathrm{Pois}(\mu_{\mathrm{image}})$. We choose our simulated lensed images to be $64\times 64$ pixels in size; see App. A for more details on the lensing forward model. The per-image likelihood is intractable because of the high dimensionality of the latent space.
@@
We would like to infer the posterior over the local as well as global parameters $p(\theta,z\mid\{x\})$ given an ensemble of images. As discussed in Sec. 2, building per-image posterior estimators ignores the joint influence of the global parameters over the entire dataset, potentially reducing the identifiability of the local parameters. If a combined posterior for the global parameters is desired, combining per-image estimators into a global posterior can also be error-prone as per-observation errors accumulate, especially if flexible density estimators (e.g., normalizing flows) are used, as we do here.
We train a hierarchical neural posterior estimator with datasets containing up to 25 lensed images, using a ResNet-18 (He et al., 2016) CNN as the per-event feature extractor. App. A describes further details related to model training.
Results
@@
Figure 6 shows exemplary posterior inference results on one of the local parameters, the Einstein radius $\theta_{E}$ (left) and one of the global parameters, the amplitude of the clump mass function $\alpha$ (right), on test datasets. The red lines show posteriors using the hierarchical estimator, while the blue dotted lines show results of using a per-image estimator. We can see that the posterior mass concentrates around the true value (vertical dashed) as more images are analyzed. On the other hand, the local parameter values are very similar to those obtained without simultaneously constraining dataset-wide global parameters. This is because, in this case, sub-clumps have a minimal (sub-percent level) effect on the lensing image, and the degeneracy between the local and global parameters is weak.
@@
Figure 6: Example results of inference on local (left) as well as global (right) gravitational lensing parameters using the hierarchical simulation-based inference model, with the posterior on global parameter shown for various cardinalities as up to 25 images are analyzed. Concentration of the posterior mass around the true global parameter value (vertical dashed lines) is seen as more lensed images are analyzed, and the posterior distributions on the local parameters are seen to be qualitatively consistent with the true parameter values used for simulation.
Statistical coverage
We perform a test of posterior statistical coverage (described in detail in Hermans et al. (2021)) for the gravitational lensing example, serving as a validation check of the learned posterior estimator across cardinalities. Using the same notation as in the main text, let $\theta$ be the global parameter of interest and $\{x\}$ be a dataset. Denoting our learned posterior density estimator as $\hat{p}(\theta|\{x\})$. For a confidence level $1-\alpha$, the expected coverage probability is
$\mathbb{E}_{(\theta,\{x\})\sim p(\theta,\{x\})}\left[\mathds{1}_{\Theta}\left(\theta\in\Theta_{\hat{p}(\theta\mid\{x\})}(1-\alpha)\right)\right],$ (13)
where $\Theta_{{\hat{p}(\theta|\{x\})}}(1-\alpha)$ gives the $1-\alpha$ highest posterior density interval (HDPI) of the estimator $\hat{p}(\theta|\{x\})$ and $\mathds{1}_{\Theta}()$ is an indicator function mapping samples that fall within the HDPI to one. Given $N$ sampled datasets from the joint distribution $(\theta^{\star},\{x\})\sim p(\theta,x)$, the empirical expected coverage for the posterior estimator $\hat{p}(\theta|\{x\})$ is
$\frac{1}{N}\sum_{i=1}^{N}\mathds{1}_{\Theta}\left(\theta^{\star}\in\Theta_{\hat{p}(\theta|\{x\})}(1-\alpha)\right).$ (14)
The nominal expected coverage is the expected coverage when $\hat{p}(\theta|\{x\})=p(\theta|\{x\})$ and is equal to the confidence level $1-\alpha$. In Fig. 7 we show the results of a test of statistical coverage for the gravitational lensing example, showing empirical against nominal coverage probability for the same posterior estimator evaluated over different cardinalities (corresponding to the different lines shown), using 100 test samples. If an estimator produces perfectly calibrated posteriors, its empirical expected coverage probability and nominal expected coverage probability match (dashed diagonal line). The amortized posteriors obtained are shown to be well-calibrated across different cardinalities.
Figure 7: Test of statistical coverage for the gravitational lensing example, showing empirical against nominal coverage probability. If an estimator produces perfectly calibrated posteriors, its empirical expected coverage probability and nominal expected coverage probability match (dashed diagonal line). The amortized posteriors obtained are shown to be well-calibrated across different cardinalities (corresponding to the different lines).
@@ 6 Conclusions
We studied the problem of simulation-based inference over event ensembles in the presence of hierarchical structures in the underlying forward model. In this setting, common in the physical sciences, hierarchy-aware parameter inference requires a dataset-wide approach. For this purpose, we introduced neural estimators for likelihood-ratios and posteriors, illustrated their use for Bayesian as well as frequentist analysis in situations typical of astro- and particle physics, and substantiated correct inference performance in these scenarios. Our methods do not require explicit parameterization in terms of marginalized or profiled nuisance parameters, making them complementary to existing techniques for dataset-wide inference. Once trained, they provide fully amortized inference over datasets of varying cardinality and offer significant speedups over traditional frequentist and Bayesian methods, making them suitable also for high-throughput applications.
@@
Limitations
Our methods apply to situations where the cardinality $N$ of the observed dataset follows an arbitrary, but fixed, distribution $p(N\mid\theta)$ (cf. equation 2). It is currently an open question to what extent estimators trained on $p_{1}(N\mid\theta)$ achieve statistically sound performance when applied to datasets corresponding to a different distribution $p_{2}(N\mid\theta)$, encoding, for example, the accumulation of additional data. We leave this extension for future investigation, but note that insights from recent work Berman et al. (2022); Chen et al. (2019) describing Bayesian updating as a dynamical system may offer a promising way to address this important problem.
Acknowledgments
We made extensive use of the Einops (Rogozhnikov, 2022), Jax (Bradbury et al., 2018), Jupyter (Kluyver et al., 2016), Matplotlib (Hunter, 2007), nflows (Durkan et al., 2020), Numpy (Harris et al., 2020), PyMC5 (Salvatier et al., 2016), PyTorch (Paszke et al., 2019), PyTorch Lightning (Falcon et al., 2020), and Scipy (Virtanen et al., 2020) packages.